# Near-linear Time Dispersion of Mobile Agents

### v2

@@ 3 Rooted Dispersion
In this section, we present an algorithm, $\mathbf{RootedDispersion}$, that solves the dispersion problem in the rooted setting. That is, it operates under the assumption that all agents are initially located at a single node $s\in V$. This algorithm straightforwardly implements the strategy of the HEO-DFS, which we presented in Section 1. The time and space complexities of this algorithm are $O(k\log\tau)$ steps and $O(\log\delta)$ bits, respectively.
In an execution of Algorithm $\mathbf{RootedDispersion}$, the agent with the largest ID, denoted as $a_{\mathrm{max}}$, serves as the leader. Note that every agent can easily determine whether it is $a_{\mathrm{max}}$ or not at time step 00 by comparing the IDs of all agents. Then, $a_{\mathrm{max}}$ conducts a depth-first search (DFS), while the other agents move with the leader and one of them settles at an unsettled node when they visit it. If $a_{\mathrm{max}}$ encounters an unsettled node without any accompanying agents, $a_{\mathrm{max}}$ settles itself on that node, achieving dispersion. During a DFS, $a_{\mathrm{max}}$ must determine (i) whether there is an unsettled neighbor of the current location, and (ii) if so, which neighbor is unsettled. To make this decision, all DFS-based algorithms in the literature have $a_{\mathrm{max}}$ visit neighbors one by one until it finds an unsettled node, which clearly requires $\Omega(\tau)$ steps. $\mathbf{RootedDispersion}$, in contrast, makes this decision in $O(\log\tau)$ steps with the help of the agents that have already settled on the neighbors of the current location.
Figure 1: The behavior of the agents when the leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at the center node $w$. A black circle, triangle, and rectangle represent a leader ($a_{\mathrm{max}}$), a non-leader explorer, and a settler, respectively. In every two time steps, the number of agents on $w$ excluding $\psi(w)$ doubles (i.e.$2\to 4\to 8$) until some agent detects an unsettled neighbor of $w$ in the sixth configuration. After that, $a_{\mathrm{max}}$ lets the helping settlers go back to their homes.
@@
2 $b.\mathtt{parent}\leftarrow\bot$ while $a_{\mathrm{max}}.\mathtt{settled}=\bot$ do
3       $\mathtt{Probe}()$ Let $w=\nu(a_{\mathrm{max}})$ if $\psi(w).\mathtt{next}\neq\bot$ then
             Let $u=N(w,\psi(w).\mathtt{next})$   // $u$ is an unsetteled node here
4             All explorers in $A(w)$ go to $u$ $b^{\prime}.\mathtt{settled}\leftarrow\top$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$ $b^{\prime}.\mathtt{parent}\leftarrow a_{\mathrm{max}}.\mathtt{pin}$
@@
7
8function $\mathtt{Probe}()$:
9       Let $w=\nu(a_{\mathrm{max}})$. $(\psi(w).\mathtt{next},\psi(w).\mathtt{checked})\leftarrow(\bot,-1)$ while $\psi(w).\mathtt{checked}\neq\delta_{w}-1$ do
10             Let $a_{1},a_{2},\dots,a_{x}$ be the agents in $A(w)\setminus\{\psi(w)\}$, and let $\delta^{\prime}=\min(x,\delta_{w}-1-\psi(w).\mathtt{checked})$. For each $i=1,2,\dots,\delta^{\prime}$, assign $a_{i}$ to the neighboring node $u_{i}=N(w,i+\psi(w).\mathtt{checked})$, and let $a_{i}$ make a round trip between $w$ and $u_{i}$. In other words, make $a_{i}$ move in the order $w\to u_{i}\to w$. If $a_{i}$ finds a settler at $u_{i}$, it will bring the settler $\psi(u_{i})$ back to $w$. if there exists $a_{i}$ that did not bring $\psi(u_{i})$ back to $w$ then
                   $\psi(w).\mathtt{next}\leftarrow i+\psi(w).\mathtt{checked}$   // $u_{i}$ must be unsettled
@@
Algorithm 1 $\mathbf{RootedDispersion}$
The pseudocode for Algorithm $\mathbf{RootedDispersion}$ is shown in Algorithm 1. This pseudocode consists of two parts: the main function (lines 1–12) and the function $\mathtt{Probe}()$ (lines 13–23). As mentioned in the previous section, every agent $a$ maintains a variable $a.\mathtt{settled}\in\{\bot,\top\}$, which decides whether $a$ is an explorer or a settler. In addition, the settler $\psi(w)$ of a node $w$ maintains two variables, $\psi(w).\mathtt{parent},\psi(w).\mathtt{next}\in[0,\delta_{w}-1]\cup{\bot}$ for the main function. As we will see later, the following are guaranteed each time $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at node $w$:
@@
Since the number of agents is $k$, the DFS-traversal stops after $a_{\mathrm{max}}$ makes a forward move $k-1$ times. The agent $a_{\mathrm{max}}$ makes a backward move at most once from any node. Therefore, excluding the execution time of $\mathtt{Probe}()$, the execution of the main function completes in $O(k)$ time. Furthermore, the function $\mathtt{Probe}()$ is invoked at most $2(k-1)$ times, once after each forward move and once after each backward move. Since a single invocation of $\mathtt{Probe}()$ requires $O(\log\tau)$ time, the overall execution time of $\mathbf{RootedDispersion}$ can be bounded by $O(k\log\tau)$ time.
@@
Let us describe the behavior of the function $\mathtt{Probe}()$, assuming that it is invoked on node $w$ at time step $t$. Figure 1 may help the readers to understand the behavior. In the execution of $\mathtt{Probe}()$, the leader $a_{\mathrm{max}}$ employs the explorers present on $w$ and (a portion of) the settlers at $N(w)$ to search for an unsettled node in $N(w)$. We implement this process with a variable $\psi(w).\mathtt{checked}\in[-1,\delta_{w}-1]$ for the settler $\psi(w)$. Specifically, explorers at node $w$ verify whether the neighbors of $w$ are unsetted or not in the order of port numbers and store the most recently checked port number in $\psi(w).\mathtt{checked}$. Consequently, $\psi(w).\mathtt{checked}=\ell$ implies that the neighbors $N(w,0),N(w,1),\dots,N(w,\ell)$ are settled. Let $x=|A(w,t)\setminus\{\psi(w)\}|$, i.e., there are $x$ agents at $w$ when $\mathtt{Probe}()$ is invoked, excluding $\psi(w)$. In the first iteration of the while loop (lines 16–22), the $\min(x,\delta_{w})$ agents concurrently visit $\min(x,\delta_{w})$ neighbors and then return to $w$ (lines 17–18). This entire process takes exactly two time steps. These agents bring back all the settlers, at most one for each neighbor, they find. If there is an agent that does not find a settler, then the node visited by that agent must be unsettled. In such a case, the port used by one of these agents is stored in $\psi(w).\mathtt{next}$, and the while loop terminates (lines 20–21).333 For simplicity, we reset $\psi(w).\mathtt{checked}$ to $-1$ each time we invoke $\mathtt{Probe}()$ at $w$, so we do not use the information about which ports were already checked in the past invocation of $\mathtt{Probe}()$. As a result, the value of $\psi(w).\mathtt{next}$ computed by $\mathtt{Probe}$ does not have to be the minimum port leading to an unsettled neighbor of $w$. If all $x$ agents bring back one agent each, then there are $2x$ agents on $w$, excluding $\psi(w)$. In the second iteration of the while loop, these $2x$ agents visit the next $2x$ neighbors and search for unsettled neighbors in a similar way. As long as no unsettled neighbors are discovered, the number of agents on $w$, excluding $\psi(w)$, doubles with each iteration of the while loop. Since there are at most $\tau=\min(k,\delta)$ settled nodes in $N(w)$, after running the while loop at most $O(\log(\tau/x))=O(\log\tau)$ times, either an unsettled node will be found, or the search will be concluded without finding any unsettled nodes. In the latter case, since $\psi(w).\mathtt{next}$ is initialized to $\bot$ when $\mathtt{Probe}()$ is called (line 15), $\psi(w).\mathtt{next}=\bot$ will also be valid at the end of the while loop, allowing $a_{\mathrm{max}}$ to verify that all neighbors of $w$ are settled. After the while loop ends, the settlers brought back to $w$ return to their homes (line 23). This process of “returning to their homes” requires the agents to remember the port number leading to their home from $w$. However, we exclude this process from the pseudocode because it can be implemented in a straightforward manner, and it requires only $O(\log\delta)$ bits of each agent’s memory. In conclusion, we have the following lemma.
@@
Each agent requires $O(\log\delta)$ bits of memory to execute $\mathbf{RootedDispersion}$.
@@
In this algorithm, an agent handles several $O(\log\delta)$-bit variable, $\mathtt{next}$, $\mathtt{checked}$, $\mathtt{parent}$, as well as the port number that the settler $\psi(u)$ needs to remember in order to return to node $u$ from node $w$ at line 23 after coming at line 17. Every other variable can be stored in a constant space. Therefore, the space complexity is $O(\log\delta)$ bits. ∎
@@
In the rooted setting, algorithm $\mathbf{RootedDispersion}$ solves the dispersion problem within $O(k\log\tau)$ time using $O(\log\delta)$ bits of space per agent.
@@
As long as there is an unsettled neighbor of the current location, $a_{\mathrm{max}}$ makes a forward move to one of those nodes. If there is no such neighbor, $a_{\mathrm{max}}$ makes a backward move to the parent node of the current location. Since the graph is connected, this DFS-traversal clearly visits $k$ nodes with exactly $k-1$ forward moves and at most $k-1$ backward moves. Thus, the number of calls to $\mathtt{Probe}()$ is at most $2(k-1)$ times. By Lemma 1, the execution of $\mathbf{RootedDispersion}$ achieves dispersion within $O(k\log\tau)$ time. ∎
@@ 4.1 Overview
In this section, we present an algorithm $\mathbf{GeneralDispersion}$ that solves the dispersion problem in $O((k\log\tau)\cdot\log k)=O(k\log^{2}k)$ time, using $O(\log(k+\delta))$ bits of each agent’s memory, in the general setting. Unlike the rooted setting, the agents are deployed arbitrarily. In $\mathbf{GeneralDispersion}$, we view the agents located at the same starting node as a single group and achieve rapid dispersion by having each group perform a HEO-DFS in parallel, sometimes merging groups. We show that by employing the group merge method given by Shintaku et al. [14], say Zombie Method, we can parallelize HEO-DFS by accepting an additive factor of $\log k$ to the space complexity and a multiplicative factor of $\log k$ to the time complexity. We have made substantial modifications to the Zombie Method to avoid conflicts between the function $\mathtt{Probe}()$ of HEO-DFS and the behavior of the Zombie Method.
As defined in Section 2, agents $a$ with $a.\mathtt{settled}=\top$ are called settlers, and the other agents are called explorers. In addition, in $\mathbf{GeneralDispersion}$, we classify explorers to two classes, leaders and zombies, depending on a variable $\mathtt{leader}\in[0,\mathrm{id}_{\mathrm{max}}]$. We call an explorer $a$ a leader if $a.\mathtt{leader}=a.\mathtt{ID}$, otherwise a zombie. Each agent $a$ initially has $a.\mathtt{leader}=a.\mathtt{ID}$, so all agents are leaders at the start of an execution of $\mathbf{GeneralDispersion}$. As we will see later, a leader may become a zombie and a zombie will eventually become a settler, whereas a zombie never becomes a leader again, and a settler never becomes a leader or zombie again. Among the agents in $A(w,t)$, the set of leaders (resp., zombies, settlers) at time step $t$ is denoted by $A_{L}(w,t)$ (resp., $A_{Z}(w,t),A_{S}(w,t)$). By definition, $A(w,t)=A_{L}(w,t)\cup A_{Z}(w,t)\cup A_{S}(w,t)$.
We introduce a variable $\mathtt{level}\in\mathbb{N}$ to bound the execution time of $\mathbf{GeneralDispersion}$. We call the value of $a.\mathtt{level}$ the level of agent $a$. The level of every agent is $1$ initially. The pair $(a.\mathtt{leader},a.\mathtt{level})$ serves as the group identifier: when agent $a$ is a leader or settler, we say that $a$ belongs to a group $(a.\mathtt{leader},a.\mathtt{level})$. By definition, for any $(\ell,i)\in\mathbb{N}^{2}$, a group $(\ell,i)$ has at most one leader. We do not use a variable $a.\mathtt{leader}$ when $a$ is a zombie, that is, a zombie does not belong to any group. However, when it accompanies a leader, it joins the HEO-DFS of that leader. We define a relationship $\prec$ between any two non-zombies $a$ and $b$ using these group identifiers as follows:
@@
$V_{a}=\{w\in V\mid\exists b\in A:\psi(w)=b\land b.\mathtt{leader}=a.\mathtt{ID}\land b.\mathtt{level}=a.\mathtt{level}\}.$
Each time a leader $a$ visits an unsettled node, it settles one of the accompanying zombies (if exists), giving it $a$’s group identifier $(a.\mathtt{leader},a.\mathtt{level})$. That is, $a$ expands its territory. If a node outside $a$’s territory is detected during the probing process of HEO-DFS, that node is considered unsettled even though it belongs to the territory of another leader. As a result, $a$ may move forward to a node $u$ that is inside another leader’s territory. If $u$ belongs to the territory of a weaker group, $a$ incorporates the settler $\psi(u)$ into its own group by giving $\psi(u)$ its group identifier $(a.\mathtt{leader},a.\mathtt{level})$. If a leader $a$ encounters a stronger leader or a stronger settler during its HEO-DFS, $a$ becomes a zombie and terminates its own HEO-DFS. If there is a leader at the current location $\nu(a)$ when $a$ becomes a zombie, $a$ joins the HEO-DFS of that leader. Otherwise, the agent $a$, now a zombie, chases a stronger leader by moving through the port $\psi(v).\mathtt{next}$ at each node $v$. Unlike $\mathbf{RootedDispersion}$, a leader updates $\psi(v).\mathtt{next}$ with the most recently used port even when it makes a backward move. This ensures that $a$ catches up to a leader eventually, at which point $a$ joins the HEO-DFS led by the leader.
Unlike $\mathbf{RootedDispersion}$, a leader does not settle itself at a node in the final stage of HEO-DFS. The leader $a$ suspends the HEO-DFS if it visits an unsettled node but it has no accompanying zombies to settle at that time. A leader who has suspended the HEO-DFS due to the absence of accompanying zombies is called a waiting leader. Conversely, a leader with accompanying zombies is called an active leader. A waiting leader $a$ resumes the HEO-DFS when a zombie catches up to $a$ at $\nu(a)$. As we will see later, the execution of $\mathbf{GeneralDispersion}$ ensures that all agents eventually become either waiting leaders or settlers, each residing at a distinct node. Since the dispersion problem does not require termination detection, the agents have solved the dispersion once such a configuration is reached.

### v3

@@ 3 Rooted Dispersion
In this section, we present an algorithm, $\mathbf{RootedDisp}$, that solves the dispersion problem in the rooted setting. That is, it operates under the assumption that all agents are initially located at a single node $s\in V$. This algorithm straightforwardly implements the strategy of the HEO-DFS, which we presented in Section 1. The time and space complexities of this algorithm are $O(k\log\tau)$ steps and $O(\log(k+\Delta))$ bits, respectively.
In an execution of Algorithm $\mathbf{RootedDisp}$, the agent with the largest ID, denoted as $a_{\mathrm{max}}$, serves as the leader. Note that every agent can easily determine whether it is $a_{\mathrm{max}}$ or not at time step 00 by comparing the IDs of all agents. Then, $a_{\mathrm{max}}$ conducts a depth-first search (DFS), while the other agents move with the leader and one of them settles at an unsettled node when they visit it. If $a_{\mathrm{max}}$ encounters an unsettled node without any accompanying agents, $a_{\mathrm{max}}$ settles itself on that node, achieving dispersion. During a DFS, $a_{\mathrm{max}}$ must determine (i) whether there is an unsettled neighbor of the current location, and (ii) if so, which neighbor is unsettled. To make this decision, all DFS-based algorithms in the literature have $a_{\mathrm{max}}$ visit neighbors one by one until it finds an unsettled node, which clearly requires $\Omega(\tau)$ steps. $\mathbf{RootedDisp}$, in contrast, makes this decision in $O(\log\tau)$ steps with the help of the agents that have already settled on the neighbors of the current location.
Figure 1: The behavior of the agents when the leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at the center node $w$ in $\mathbf{RootedDisp}$. A black circle, triangle, and rectangle represent a leader ($a_{\mathrm{max}}$), a non-leader explorer, and a settler, respectively. The integers in the leftmost figure represents port numbers. In every two time steps, the number of agents on $w$ excluding $\psi(w)$ doubles (i.e.$2\to 4\to 8$) until some agent detects an unsettled neighbor of $w$. After that, $a_{\mathrm{max}}$ lets the helping settlers go back to their homes.
@@
2 $b.\mathtt{parent}\leftarrow\bot$
3 while $a_{\mathrm{max}}.\mathtt{settled}=\bot$ do
4       $\mathtt{Probe}()$
5       Let $w=\nu(a_{\mathrm{max}})$
6       if $\psi(w).\mathtt{next}\neq\bot$ then
             Let $u=N(w,\psi(w).\mathtt{next})$
              // $u$ is an unsetteled node here
7             All explorers in $A(w)$ go to $u$
8             $b^{\prime}.\mathtt{settled}\leftarrow\top$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$
9             $b^{\prime}.\mathtt{parent}\leftarrow a_{\mathrm{max}}.\mathtt{pin}$
@@ 7             All explorers in $A(w)$ go to $u$
12
13
14
15function $\mathtt{Probe}()$:
16       Let $w=\nu(a_{\mathrm{max}})$.
17       $(\psi(w).\mathtt{next},\psi(w).\mathtt{checked})\leftarrow(\bot,-1)$
18       while $\psi(w).\mathtt{checked}\neq\delta_{w}-1$ do
19             Let $a_{1},a_{2},\dots,a_{x}$ be the agents in $A(w)\setminus\{\psi(w)\}$, and let $\Delta^{\prime}=\min(x,\delta_{w}-1-\psi(w).\mathtt{checked})$. For each $i=1,2,\dots,\Delta^{\prime}$, assign $a_{i}$ to the neighboring node $u_{i}=N(w,i+\psi(w).\mathtt{checked})$, and let $a_{i}$ make a round trip between $w$ and $u_{i}$. In other words, make $a_{i}$ move in the order $w\to u_{i}\to w$. If $a_{i}$ finds a settler at $u_{i}$, it will bring the settler $\psi(u_{i})$ back to $w$.
20             if there exists $a_{i}$ that did not bring $\psi(u_{i})$ back to $w$ then
                   $\psi(w).\mathtt{next}\leftarrow i+\psi(w).\mathtt{checked}$
                    // $u_{i}$ must be unsettled
@@
Algorithm 1 $\mathbf{RootedDisp}$
The pseudocode for Algorithm $\mathbf{RootedDisp}$ is shown in Algorithm 1. This pseudocode consists of two parts: the main function (lines 1–12) and the function $\mathtt{Probe}()$ (lines 13–23). As mentioned in the previous section, every agent $a$ maintains a variable $a.\mathtt{settled}\in\{\bot,\top\}$, which decides whether $a$ is an explorer or a settler. In addition, the settler $\psi(w)$ of a node $w$ maintains two variables, $\psi(w).\mathtt{parent},\psi(w).\mathtt{next}\in[0,\delta_{w}-1]\cup{\bot}$ for the main function. As we will see later, the following are guaranteed each time $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at node $w$:
@@
Since the number of agents is $k$, the DFS-traversal stops after $a_{\mathrm{max}}$ makes a forward move $k-1$ times. The agent $a_{\mathrm{max}}$ makes a backward move at most once from any node. Therefore, excluding the execution time of $\mathtt{Probe}()$, the execution of the main function completes in $O(k)$ time. Furthermore, the function $\mathtt{Probe}()$ is invoked at most $2(k-1)$ times, once after each forward move and once after each backward move. Since a single invocation of $\mathtt{Probe}()$ requires $O(\log\tau)$ time, the overall execution time of $\mathbf{RootedDisp}$ can be bounded by $O(k\log\tau)$ time.
@@
Let us describe the behavior of the function $\mathtt{Probe}()$, assuming that it is invoked on node $w$ at time step $t$. Figure 1 may help the readers to understand the behavior. In the execution of $\mathtt{Probe}()$, the leader $a_{\mathrm{max}}$ employs the explorers present on $w$ and (a portion of) the settlers at $N(w)$ to search for an unsettled node in $N(w)$. We implement this process with a variable $\psi(w).\mathtt{checked}\in[-1,\delta_{w}-1]$ for the settler $\psi(w)$. Specifically, explorers at node $w$ verify whether the neighbors of $w$ are unsettled or not in the order of port numbers and store the most recently checked port number in $\psi(w).\mathtt{checked}$. Consequently, $\psi(w).\mathtt{checked}=\ell$ implies that the neighbors $N(w,0),N(w,1),\dots,N(w,\ell)$ are settled. Let $x=|A(w,t)\setminus\{\psi(w)\}|$, i.e., there are $x$ agents at $w$ when $\mathtt{Probe}()$ is invoked, excluding $\psi(w)$. In the first iteration of the while loop (lines 16–22), the $\min(x,\delta_{w})$ agents concurrently visit $\min(x,\delta_{w})$ neighbors and then return to $w$ (lines 17–18). This entire process takes exactly two time steps. These agents bring back all the settlers, at most one for each neighbor, they find. If there is an agent that does not find a settler, then the node visited by that agent must be unsettled. In such a case, the port used by one of these agents is stored in $\psi(w).\mathtt{next}$, and the while loop terminates (lines 20–21).222 For simplicity, we reset $\psi(w).\mathtt{checked}$ to $-1$ each time we invoke $\mathtt{Probe}()$ at $w$, so we do not use the information about which ports were already checked in the past invocation of $\mathtt{Probe}()$. As a result, the value of $\psi(w).\mathtt{next}$ computed by $\mathtt{Probe}$ does not have to be the minimum port leading to an unsettled neighbor of $w$. If all $x$ agents bring back one agent each, then there are $2x$ agents on $w$, excluding $\psi(w)$. In the second iteration of the while loop, these $2x$ agents visit the next $2x$ neighbors and search for unsettled neighbors in a similar way. As long as no unsettled neighbors are discovered, the number of agents on $w$, excluding $\psi(w)$, doubles with each iteration of the while loop. Since there are at most $\tau=\min(k,\Delta)$ settled nodes in $N(w)$, after running the while loop at most $O(\log(\tau/x))=O(\log\tau)$ times, either an unsettled node will be found, or the search will be concluded without finding any unsettled nodes. In the latter case, since $\psi(w).\mathtt{next}$ is initialized to $\bot$ when $\mathtt{Probe}()$ is called (line 15), $\psi(w).\mathtt{next}=\bot$ will also be valid at the end of the while loop, allowing $a_{\mathrm{max}}$ to verify that all neighbors of $w$ are settled. After the while loop ends, the settlers brought back to $w$ return to their homes (line 23). This process of “returning to their homes” requires the agents to remember the port number leading to their home from $w$. However, we exclude this process from the pseudocode because it can be implemented in a straightforward manner, and it requires only $O(\log\Delta)$ bits of each agent’s memory. In conclusion, we have the following lemma.
@@
Each agent requires $O(\log(k+\Delta))$ bits of memory to execute $\mathbf{RootedDisp}$.
@@
In this algorithm, an agent handles several $O(\log\Delta)$-bit variable, $\mathtt{next}$, $\mathtt{checked}$, $\mathtt{parent}$, as well as the port number that the settler $\psi(u)$ needs to remember in order to return to node $u$ from node $w$ at line 23 after coming at line 17. Every other variable can be stored in a constant space. Therefore, the space complexity is $O(\log(k+\Delta))$ bits, adding the memory space to store the agent’s identifier. ∎
@@
In the rooted setting, algorithm $\mathbf{RootedDisp}$ solves the dispersion problem within $O(k\log\tau)$ time using $O(\log(k+\Delta))$ bits of space per agent.
@@
As long as there is an unsettled neighbor of the current location, $a_{\mathrm{max}}$ makes a forward move to one of those nodes. If there is no such neighbor, $a_{\mathrm{max}}$ makes a backward move to the parent node of the current location. Since the graph is connected, this DFS-traversal clearly visits $k$ nodes with exactly $k-1$ forward moves and at most $k-1$ backward moves. Thus, the number of calls to $\mathtt{Probe}()$ is at most $2(k-1)$ times. By Lemma 1, the execution of $\mathbf{RootedDisp}$ achieves dispersion within $O(k\log\tau)$ time. ∎
@@ 4.1 Overview
In this section, we present an algorithm $\mathbf{GeneralDisp}$ that solves the dispersion problem in $O(k\log\tau\cdot\log k)=O(k\log^{2}k)$ time, using $O(\log(k+\Delta))$ bits of each agent’s memory, in the general setting. Unlike the rooted setting, the agents are deployed arbitrarily. In $\mathbf{GeneralDisp}$, we view the agents located at the same starting node as a single group and achieve rapid dispersion by having each group perform a HEO-DFS in parallel, sometimes merging groups. We show that by employing the group merge method given by Shintaku et al. [16], say Zombie Method, we can parallelize HEO-DFS by accepting an additive factor of $\log k$ to the space complexity and a multiplicative factor of $\log k$ to the time complexity. We have made substantial modifications to the Zombie Method to avoid conflicts between the function $\mathtt{Probe}()$ of HEO-DFS and the behavior of the Zombie Method.
As defined in Section 2, agents $a$ with $a.\mathtt{settled}=\top$ are called settlers, and the other agents are called explorers. In addition, in $\mathbf{GeneralDisp}$, we classify explorers to two classes, leaders and zombies, depending on a variable $\mathtt{leader}\in[0,\mathrm{id}_{\mathrm{max}}]$. We call an explorer $a$ a leader if $a.\mathtt{leader}=a.\mathtt{ID}$, otherwise a zombie. Each agent $a$ initially has $a.\mathtt{leader}=a.\mathtt{ID}$, so all agents are leaders at the start of an execution of $\mathbf{GeneralDisp}$. As we will see later, a leader may become a zombie and a zombie will eventually become a settler, whereas a zombie never becomes a leader again, and a settler never becomes a leader or zombie again. Among the agents in $A(v,t)$, the set of leaders (resp., zombies, settlers) staying at $v$ in time step $t$ is denoted by $A_{L}(v,t)$ (resp., $A_{Z}(v,t),A_{S}(v,t)$). By definition, $A(v,t)=A_{L}(v,t)\cup A_{Z}(v,t)\cup A_{S}(v,t)$.
We introduce a variable $\mathtt{level}\in\mathbb{N}$ to bound the execution time of $\mathbf{GeneralDisp}$. We call the value of $a.\mathtt{level}$ the level of agent $a$. The level of every agent is $1$ initially. The pair $(a.\mathtt{leader},a.\mathtt{level})$ serves as the group identifier: when agent $a$ is a leader or settler, we say that $a$ belongs to a group $(a.\mathtt{leader},a.\mathtt{level})$. By definition, for any $(\ell,i)\in\mathbb{N}^{2}$, a group $(\ell,i)$ has at most one leader. A zombie does not belong to any group. However, when it accompanies a leader, it joins the HEO-DFS of that leader. We define a relationship $\prec$ between any two non-zombies $a$ and $b$ using these group identifiers as follows:
@@
$V_{a}=\{v\in V\mid\exists b\in A:\psi(v)=b\land b.\mathtt{leader}=a.\mathtt{ID}\land b.\mathtt{level}=a.\mathtt{level}\}.$
Each time a leader $a$ visits an unsettled node, it settles one of the accompanying zombies (if exists), giving it $a$’s group identifier $(a.\mathtt{leader},a.\mathtt{level})$. That is, $a$ expands its territory. If a node outside $a$’s territory is detected during the probing process of HEO-DFS, that node is considered unsettled even though it belongs to the territory of another leader. As a result, $a$ may move forward to a node $u$ that is inside another leader’s territory. If that node $u$ belongs to the territory of a weaker group, $a$ incorporates the settler $\psi(u)$ into its own group by giving $\psi(u)$ its group identifier $(a.\mathtt{leader},a.\mathtt{level})$. If a leader $a$ encounters a stronger leader or a stronger settler during its HEO-DFS, $a$ becomes a zombie and terminates its own HEO-DFS. If there is a leader at the current location $\nu(a)$ when $a$ becomes a zombie, $a$ joins the HEO-DFS of that leader. Otherwise, the agent $a$, now a zombie, chases a stronger leader by moving through the port $\psi(v).\mathtt{next}$ at each node $v$. Unlike $\mathbf{RootedDisp}$, a leader updates $\psi(v).\mathtt{next}$ with the most recently used port even when it makes a backward move. This ensures that $a$ catches up to a leader eventually, at which point $a$ joins the HEO-DFS led by the leader.
Unlike $\mathbf{RootedDisp}$, a leader does not settle itself at a node in the final stage of HEO-DFS. The leader $a$ suspends the HEO-DFS if it visits an unsettled node but it has no accompanying zombies to settle at that time. A leader who has suspended the HEO-DFS due to the absence of accompanying zombies is called a waiting leader. Conversely, a leader with accompanying zombies is called an active leader. A waiting leader $a$ resumes the HEO-DFS when a zombie catches up to $a$ at $\nu(a)$. As we will see later, the execution of $\mathbf{GeneralDisp}$ ensures that all agents eventually become either waiting leaders or settlers, each residing at a distinct node. The agents have solved the dispersion once such a configuration is reached because thereafter no agent moves and no two agents are co-located.